OpenStudy (anonymous):

HELP PLEASE!!! Given the parent functions f(x) = log3 (5x − 5) and g(x) = log3 (x − 1), what is f(x) − g(x)? f(x) − g(x) = log3 (4x − 4) f(x) − g(x) = log3 (4x − 6) f(x) − g(x) = log3 5 f(x) − g(x) = = log3 1/5

OpenStudy (anonymous):

f(x) - g(x) = log3 (5x - 5) - log3 (x - 1) = log3 (5x - 5) / (x - 1) = log3 5(x - 1) / (x - 1) = log3 5

OpenStudy (anonymous):

yeah i agree with @Siri15

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 3x − 1, what is f(x) • g(x)? f(x) • g(x) = log10 (3x − 1)^x f(x) • g(x) = log10 x^(3x − 1) f(x) • g(x) = 3x log10 x + log10 x f(x) • g(x) = log10 x − 3x log10 x